The cash records of Pharoah Company show the following.

For July:

1. The June 30 bank reconciliation indicated that deposits in transit total $730. During July, the general ledger account Cash shows deposits of $17,020, but the bank statement indicates that only $15,720 in deposits were received during the month.
2. The June 30 bank reconciliation also reported outstanding checks of $970. During the month of July, Pharoah Company books show that $19,420 of checks were issued, yet the bank statement showed that $16,350 of checks cleared the bank in July.


For September:

3. In September, deposits per bank statement totaled $26,770, deposits per books were $25,840, and deposits in transit at September 30 were $2,870.
4. In September, cash disbursements per books were $23,440, checks clearing the bank were $24,570, and outstanding checks at September 30 were $2,440.


There were no bank debit or credit memoranda, and no errors were made by either the bank or Pharoah Company.

Answer the following questions.

(a) In situation 1, what were the deposits in transit at July 31?

The deposits in transit at July 31 $enter the deposits in transit at July 31 in dollars


(b) In situation 2, what were the outstanding checks at July 31?

The outstanding checks at July 31 $enter the outstanding checks at July 31 in dollars


(c) In situation 3, what were the deposits in transit at August 31?

The deposits in transit at August 31 $enter the deposits in transit at August 31 in dollars


(d) In situation 4, what were the outstanding checks at August 31?

The outstanding checks at August 31 $enter the outstanding checks at August 31 in dollars
Answer #1

a) The deposit in transit at July 31 = 730+17020-15720 = 2030

b) Outstanding Checks at July 31 = 970+19420-16350 = 4040

c) deposit in transit at august 31 = 2870+26770-25840 = 3800

d) Outstanding checks at August 31 = 2440+24570-23440 = 3570